# MEMOIR: Lifelong Model Editing with Minimal Overwrite and Informed Retention for LLMs

### v3

@@ Abstract
Language models deployed in real-world systems often require post-hoc updates to incorporate new or corrected knowledge. However, editing such models efficiently and reliably—without retraining or forgetting previous information—remains a major challenge. Existing methods for lifelong model editing either compromise generalization, interfere with past edits, or fail to scale to long editing sequences. We propose MEMOIR, a novel scalable framework that injects knowledge through a residual memory, i.e., a dedicated parameter module, while preserving the core capabilities of the pre-trained model. By sparsifying input activations through sample-dependent masks, MEMOIR confines each edit to a distinct subset of the memory parameters, minimizing interference among edits. At inference, it identifies relevant edits by comparing the sparse activation patterns of new queries to those stored during editing. This enables generalization to rephrased queries by activating only the relevant knowledge while suppressing unnecessary memory activation for unrelated prompts. Experiments on question answering, hallucination correction, and out-of-distribution generalization benchmarks for LLaMA-3 and Mistral backbones demonstrate that MEMOIR achieves state-of-the-art performance across reliability, generalization, and locality metrics, scaling to thousands of sequential edits with minimal forgetting.
@@ 1 Introduction
Large language models (LLMs) exhibit strong performance across a wide range of tasks, enabled by large-scale pre-training on diverse and extensive datasets (gpt-3, ; Chowdhery2022PaLMSL, ). However, LLMs can generate outdated or inaccurate information and can perpetuate biases during deployment llm_factual_error ; hallucination_survey ; chatgpt_bias , which requires further continuous updates to their knowledge base after pre-training to correct these behaviors. Although fine-tuning on new corpora can revise model behavior, it is computationally expensive and prone to catastrophic forgetting  (Goodfellow2013AnEI, ; Oswald2019ContinualLW, ). These limitations motivate an emerging line of work known as lifelong model editing (grace, ; wang2024wise, ), which aims to continuously update the model’s knowledge in an efficient and localized manner.
@@
Previous methods for knowledge editing can be broadly categorized into non-parametric and parametric approaches. Non-parametric methods (grace, ; zhang2025resolving, ) store fixed input-output activation patterns to directly correct the model’s predictions during inference. While this enables precise and localized knowledge edits, such rigid memorization typically generalizes poorly to semantically similar queries. In contrast, parametric approaches (rome, ; serac, ; t-patcher, ; wang2024wise, ; memit, ; fang2025alphaedit, ) modify a subset of model parameters—either within the original network or in a dedicated memory module—to integrate new knowledge. Although these methods offer better generalization to related inputs, continuous editing can still cause newer updates to overwrite earlier ones, resulting in catastrophic forgetting of previous edited knowledge and limiting their reliability over long sequences of edits.
Figure 1: Left: trade-offs among reliability (Rel.), generalization (Gen.), and locality (Loc.) for 1,0001{,}000 continual edits. Right: Average performance of Rel., Gen., and Loc. under varying numbers of edits. Both evaluated on LLaMA-3-8B-Instruct (llama-3, ) with ZsRE dataset (zsre, ). MEMOIR delivers the best balance among the three metrics and scales with a large number of edits.
In this paper, we propose MEMOIR, (Model Editing with Minimal Overwrite and Informed Retention), which delivers a desirable balance between reliability, generalization, and locality for a large number of edits (Figure 1). MEMOIR introduces a memory module, i.e., a dedicated fully-connected layer in a single transformer block where all edits are performed. MEMOIR mitigates catastrophic forgetting by allocating distinct parameter subsets to each edit and retrieving them during inference to activate only the relevant knowledge for a given prompt. During editing, we propose a structured sparsification of the input with sample-dependent masks to this layer, dynamically activating only a sample-specific subset of parameters in the introduced memory module in the forward pass. MEMOIR therefore promotes the distribution of new knowledge across the parameter space, reducing the overwriting of previous updates and significantly mitigating catastrophic forgetting. During inference, we use the sparsification pattern of a prompt to infer if it semantically corresponds to an edited prompt, and route the activations accordingly. This targeted knowledge activation eliminates the need for large corpora of irrelevant samples to regularize training (memit, ; fang2025alphaedit, ; wang2024wise, ). Specifically, if the input corresponds to a rephrased version of an edit, we activate only the relevant knowledge of that edit to align the representation of the prompt. Conversely, the introduced module is deactivated when detecting irrelevant prompts, effectively preserving the pre-trained knowledge of the LLM.
@@
  • We extensively evaluate MEMOIR on Q&A, hallucination correction, OOD generalization, and multi-hop reasoning tasks, demonstrating state-of-the-art results across LLaMA-3 (llama-3, ), Mistral (mistral, ), LLaMA-2 (llama-2, ) and GPT-J (gpt-j, ) architectures compared to all previous methods.
@@ 2 Lifelong Model Editing
Lifelong model editing aims to continuously add new knowledge into a model without forgetting previously acquired knowledge (grace, ; wang2024wise, ). Let $\bm{\theta}\in\mathbb{R}^{d}$ be the parameters of an LLM denoted as a function $f_{\bm{\theta}}:\mathcal{X}\to\mathcal{Y}$. It maps an input prompt $\bm{x}\in\mathcal{X}$ to an output prediction $f_{\bm{\theta}}(\bm{x})\in\mathcal{Y}$. The LLM has been trained with internet-scale corpora $\mathcal{D}_{\textrm{train}}$. During editing, the model receives a time-evolving stream of edit samples $\mathcal{D}_{\textrm{edit}}=\{(\bm{x}_{e}^{t},\bm{y}_{e}^{t})\}_{t}$, where each pair $(\bm{x}_{e}^{t},\bm{y}_{e}^{t})\in\mathcal{D}_{\text{edit}}$ represents the $t$-th edit sample in the edit sequence. In real-world settings, this sequence is expected to grow continually, potentially reaching thousands of edits.
@@
LLMs are typically based on the transformer architecture vaswani2017attention and consist of $L$ identical blocks containing multi-head attention modules followed by feed-forward network (FFN) modules. For each block $\ell\in[L]$, the FFN module consists of two fully connected layers, parameterized by $\bm{W}_{\text{fc}}^{\ell}$ and $\bm{W}_{\text{proj}}^{\ell}$, and an activation function. Previous work shows that knowledge is typically stored in the middle blocks of an LLM rome , and updating the parameters of a single middle block rather than fine-tuning the entire model can be effective to edit factual knowledge wang2024wise ; grace . In what follows, to lighten notation, we will omit the layer-indexing superscript. Let $\bm{h}$ represent the output from the attention module; then the output of the FFN module is formally defined as:
@@
for an element-wise activation function $\sigma(\cdot)$. The FFN module can be interpreted as a two-layer key-value memory system (geva-etal-2021-transformer, ), where the second layer $\bm{W}_{\text{proj}}$ functions as a linear associative memory (rome, ), mapping a sequence of input vectors $\bm{a}_{1},\bm{a}_{2},\dots]$ to corresponding output vectors $\bm{v}_{1},\bm{v}_{2},\dots]$. This memory-like behavior enables information retrieval, where $\bm{a}=\sigma(\bm{W}_{\text{fc}}\ \bm{h})$ and $\bm{v}=\text{FFN}(\bm{h})$ respectively represent the input and output vectors of the layer $\bm{W}_{\text{proj}}$. Leveraging this perspective, previous works have shown the effectiveness of modifying the output of $\bm{W}_{\text{proj}}$ to update the model’s stored knowledge (rome, ; geva-etal-2021-transformer, ; wang2024wise, ; memit, ; fang2025alphaedit, ). However, continuously updating $\bm{W}_{\text{proj}}$ in lifelong editing gradually overwrites previous edits, leading to significant forgetting and severe performance degradation with a large number of edits, as shown in Figure 1.
@@ 3 Memory Editing with Minimal Overwrite and Informed Retention
Following prior works rome ; fang2025alphaedit , we solely modify the output of a projection layer of the FFN $\bm{W}_{\text{proj}}$ of a specific block $\ell$ , which is denoted as $\bm{W}_{0}$ in the following. To preserve the knowledge stored in the pre-trained model, we do not directly modify the pre-trained weights of this layer but introduce a residual memory layer $\bm{W}_{\text{m}}$ to incorporate new edits. $\bm{W}_{\text{m}}$ is a zero-initialized copy of the original matrix $\bm{W}_{0}$ such that it contains no information before editing. The final output of the edited layer combines the output of the original layer and the residual memory layer.
@@ 3.1 Editing: distributing knowledge across memory
Prior parametric approaches on knowledge editing memit ; fang2025alphaedit typically store the knowledge of each edit across the entire trainable parameter space. However, unconstrained updates on all trainable parameters lead to severe catastrophic forgetting, as new edits quickly overwrite knowledge from previous ones, as shown in Figure 1. In this work, we draw motivation from the continual learning literature leveraging sparsity mallya2018packnet ; wortsman2020supermasks ; mirzadeh2020understanding ; saha2021gradient ; wang2022sparcl ; abbasi2022sparsity to mitigate catastrophic forgetting, and propose a structured sparsification of input activations to the residual memory during the forward pass. This ensures that each edit modifies only a subset of the trainable parameters, promoting less overlap in representations and leading to more localized updates across different edits.
@@
A fundamental aspect of this sparse masking is the selection mechanism for the active indices. Intuitively, this mechanism should satisfy two different criteria: selection of feature diversity and semantic coherence. Diversity in feature selection encourages the model to consider a broader set of features, reducing the risk of overfitting on dominant features and mitigating catastrophic forgetting by distributing updates across the trainable parameters of the memory module. Semantic coherence ensures that semantically similar input prompts yield similar selected masks, facilitating the activation of relevant knowledge in the memory module for previously unseen yet semantically related queries. Therefore, we opt for structured sparsity meier2008group .
@@
defeating the purpose of structured and reusable knowledge storage. To address this limitation, we turn to a more robust fingerprinting mechanism based on the inherent structure of the activations themselves. For an input $\bm{a}=\bm{a}(\bm{x})\in\mathbb{R}^{D}$, we compute a top-k selection based on feature magnitudes, mapping $\bm{a}$ to a binary mask $\mathcal{T}(\bm{a})$. Denoting as $a_{(k)}$ the $k$-th largest value using order statistics notation, the mask is formally defined as $\mathcal{T}(\bm{a})_{j}=\mathds{1}[a_{j}\geq a_{(k)}],\;j\in[D]$. This strategy ensures that semantically similar input prompts—naturally producing similar activation patterns—yield similar masks, directly addressing the generalization shortcoming of random seeds. Given the structured and highly correlated nature of activations in LLMs, this step effectively acts as a Locality-Sensitive Hashing (LSH) indyk1998approximate mechanism, mapping semantically similar input prompts to identical or closely related binary codes norouzi2012hamming .
@@ 3.2 Inference: activating knowledge conditional to inference prompt
During editing, we construct a database of sparse masks of each edited sample. Let $\bm{x}\in\mathcal{D}_{\textrm{edit}}$ be an edited prompt. Given that the layers prior to the introduced memory module are frozen, the application of the TopHash mechanism will yield the same sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ during both editing and inference, allowing the reliable identification of the prompt as part of the edit set by finding the exact match in the above database. Since transformers cluster semantically similar input prompts together due to their ability to generalize across paraphrases sentencebert , we expect that $\bm{x}$ should have activations, as well as TopHash masks, with high similarity w.r.t. its semantically rephrased variants. Conversely, semantically unrelated prompts will exhibit low similarity in both activation and mask space to the edited prompt $\bm{x}$.
We can compare the semantic similarity among prompts via the Hamming distance, denoted as $d_{H}$, between their respective masks norouzi2012hamming . By populating the database during editing, we can identify the semantically closest edited prompt to the prompt $\bm{x}$:
@@
The overlapping ratio $R_{\textrm{match}}$ measures the semantic relevance between a prompt and its closest edited prompt, enabling us to distinguish between different prompt types. Figure 4 presents the distribution of $R_{\textrm{match}}$ for edited prompts, rephrased variants, and irrelevant prompts across two benchmarks. We observe that these prompt categories exhibit distinct modes within the distribution, demonstrating the effectiveness of the proposed TopHash mechanism as a Locality-Sensitive Hashing (LSH) function indyk1998approximate that clusters semantically similar input prompts.
@@
Prior approaches often require access to large corpora of irrelevant samples to enhance locality; either by sampling a different irrelevant sample during each edit (wang2024wise, ) or by performing a forward pass on a corpora of up to 100,000100{,}000 samples to regularize editing (memit, ; fang2025alphaedit, ). MEMOIR, on the other hand, eliminates the need for such large-scale irrelevant samples during editing and offers a more flexible alternative, as it relies on the introduced conditional knowledge activation mechanism in Equation 4 to decide whether to use the edited or pre-trained knowledge.

### v4

@@ Abstract
Language models deployed in real-world systems often require post-hoc updates to incorporate new or corrected knowledge. However, editing such models efficiently and reliably—without retraining or forgetting previous information—remains a major challenge. Existing methods for lifelong model editing either compromise generalization, interfere with past edits, or fail to scale to long editing sequences. We propose MEMOIR, a novel scalable framework that injects knowledge through a residual memory, i.e., a dedicated parameter module, while preserving the core capabilities of the pre-trained model. By sparsifying input activations through sample-dependent masks, MEMOIR confines each edit to a distinct subset of the memory parameters, minimizing interference among edits. At inference, it identifies relevant edits by comparing the sparse activation patterns of new queries to those stored during editing. This enables generalization to rephrased queries by activating only the relevant knowledge while suppressing unnecessary memory activation for unrelated prompts. Experiments on question answering, hallucination correction, and out-of-distribution generalization benchmarks for LLaMA-3 and Mistral backbones demonstrate that MEMOIR achieves state-of-the-art performance across reliability, generalization, and locality metrics, scaling to thousands of sequential edits with minimal forgetting.  [^1] Code at https://github.com/qym7/MEMOIR
@@ 1 Introduction
Large language models (LLMs) exhibit strong performance across a wide range of tasks, enabled by large-scale pre-training on diverse and extensive datasets (Brown et al., 2020; Chowdhery et al., 2022). However, LLMs can generate outdated or inaccurate information and can perpetuate biases during deployment Balachandran et al. (2022); Ji et al. (2023); Ferrara (2023), which requires further continuous updates to their knowledge base after pre-training to correct these behaviors. Although fine-tuning on new corpora can revise model behavior, it is computationally expensive and prone to catastrophic forgetting  (Goodfellow et al., 2013; von Oswald et al., 2019). These limitations motivate an emerging line of work known as lifelong model editing (Hartvigsen et al., 2023; Wang et al., 2024), which aims to continuously update the model’s knowledge in an efficient and localized manner.
@@
Previous methods for knowledge editing can be broadly categorized into non-parametric and parametric approaches. Non-parametric methods (Hartvigsen et al., 2023; Zhang et al., 2025) store fixed input-output activation patterns to directly correct the model’s predictions during inference. While this enables precise and localized knowledge edits, such rigid memorization typically generalizes poorly to semantically similar queries. In contrast, parametric approaches (Meng et al., 2022; Mitchell et al., 2022b; Huang et al., 2023; Wang et al., 2024; Meng et al., 2023; Fang et al., 2025) modify a subset of model parameters—either within the original network or in a dedicated memory module—to integrate new knowledge. Although these methods offer better generalization to related inputs, continuous editing can still cause newer updates to overwrite earlier ones, resulting in catastrophic forgetting of previous edited knowledge and limiting their reliability over long sequences of edits.
Figure 1: Left: trade-offs among reliability (Rel.), generalization (Gen.), and locality (Loc.) for 1,0001{,}000 continual edits. Right: Average performance of Rel., Gen., and Loc. under varying numbers of edits. Both evaluated on LLaMA-3-8B-Instruct (Aaron et al., 2024) with ZsRE dataset (Levy et al., 2017). MEMOIR delivers the best balance among the three metrics and scales with a large number of edits.
In this paper, we propose MEMOIR, (Model Editing with Minimal Overwrite and Informed Retention), which delivers a desirable balance between reliability, generalization, and locality for a large number of edits (Figure 1). MEMOIR introduces a memory module, i.e., a dedicated fully-connected layer in a single transformer block where all edits are performed. MEMOIR mitigates catastrophic forgetting by allocating distinct parameter subsets to each edit and retrieving them during inference to activate only the relevant knowledge for a given prompt. During editing, we propose a structured sparsification of the input with sample-dependent masks to this layer, dynamically activating only a sample-specific subset of parameters in the introduced memory module in the forward pass. MEMOIR therefore promotes the distribution of new knowledge across the parameter space, reducing the overwriting of previous updates and significantly mitigating catastrophic forgetting. During inference, we use the sparsification pattern of a prompt to infer if it semantically corresponds to an edited prompt, and route the activations accordingly. This targeted knowledge activation eliminates the need for large corpora of irrelevant samples to regularize training (Meng et al., 2023; Fang et al., 2025; Wang et al., 2024). Specifically, if the input corresponds to a rephrased version of an edit, we activate only the relevant knowledge of that edit to align the representation of the prompt. Conversely, the introduced module is deactivated when detecting irrelevant prompts, effectively preserving the pre-trained knowledge of the LLM.
@@
  • We extensively evaluate MEMOIR on Q&A, hallucination correction, OOD generalization, and multi-hop reasoning tasks, demonstrating state-of-the-art results across LLaMA-3 (Aaron et al., 2024), Mistral (Jiang et al., 2023), LLaMA-2 (Touvron et al., 2023) and GPT-J (Wang and Komatsuzaki, 2021) architectures compared to all previous methods.
@@ 2 Lifelong Model Editing
Lifelong model editing aims to continuously add new knowledge into a model without forgetting previously acquired knowledge (Hartvigsen et al., 2023; Wang et al., 2024). Let $\bm{\theta}\in\mathbb{R}^{d}$ be the parameters of an LLM denoted as a function $f_{\bm{\theta}}:\mathcal{X}\to\mathcal{Y}$. It maps an input prompt $\bm{x}\in\mathcal{X}$ to an output prediction $f_{\bm{\theta}}(\bm{x})\in\mathcal{Y}$. The LLM has been trained with internet-scale corpora $\mathcal{D}_{\textrm{train}}$. During editing, the model receives a time-evolving stream of edit samples $\mathcal{D}_{\textrm{edit}}=\{(\bm{x}_{e}^{t},\bm{y}_{e}^{t})\}_{t}$, where each pair $(\bm{x}_{e}^{t},\bm{y}_{e}^{t})\in\mathcal{D}_{\text{edit}}$ represents the $t$-th edit sample in the edit sequence. In real-world settings, this sequence is expected to grow continually, potentially reaching thousands of edits.
@@
LLMs are typically based on the transformer architecture Vaswani et al. (2017) and consist of $L$ identical blocks containing multi-head attention modules followed by feed-forward network (FFN) modules. For each block $\ell\in[L]$, the FFN module consists of two fully connected layers, parameterized by $\bm{W}_{\text{fc}}^{\ell}$ and $\bm{W}_{\text{proj}}^{\ell}$, and an activation function. Previous work shows that knowledge is typically stored in the middle blocks of an LLM Meng et al. (2022), and updating the parameters of a single middle block rather than fine-tuning the entire model can be effective to edit factual knowledge Wang et al. (2024); Hartvigsen et al. (2023). In what follows, to lighten notation, we will omit the layer-indexing superscript. Let $\bm{h}$ represent the output from the attention module; then the output of the FFN module is formally defined as:
@@
for an element-wise activation function $\sigma(\cdot)$. The FFN module can be interpreted as a two-layer key-value memory system (Geva et al., 2021), where the second layer $\bm{W}_{\text{proj}}$ functions as a linear associative memory (Meng et al., 2022), mapping a sequence of input vectors $\bm{a}_{1},\bm{a}_{2},\dots]$ to corresponding output vectors $\bm{v}_{1},\bm{v}_{2},\dots]$. This memory-like behavior enables information retrieval, where $\bm{a}=\sigma(\bm{W}_{\text{fc}}\ \bm{h})$ and $\bm{v}=\text{FFN}(\bm{h})$ respectively represent the input and output vectors of the layer $\bm{W}_{\text{proj}}$. Leveraging this perspective, previous works have shown the effectiveness of modifying the output of $\bm{W}_{\text{proj}}$ to update the model’s stored knowledge (Meng et al., 2022; Geva et al., 2021; Wang et al., 2024; Meng et al., 2023; Fang et al., 2025). However, continuously updating $\bm{W}_{\text{proj}}$ in lifelong editing gradually overwrites previous edits, leading to significant forgetting and severe performance degradation with a large number of edits, as shown in Figure 1.
@@ 3 Memory Editing with Minimal Overwrite and Informed Retention
Following prior works Meng et al. (2022); Fang et al. (2025), we solely modify the output of a projection layer of the FFN $\bm{W}_{\text{proj}}$ of a specific block $\ell$ , which is denoted as $\bm{W}_{0}$ in the following. To preserve the knowledge stored in the pre-trained model, we do not directly modify the pre-trained weights of this layer but introduce a residual memory layer $\bm{W}_{\text{m}}$ to incorporate new edits. $\bm{W}_{\text{m}}$ is a zero-initialized copy of the original matrix $\bm{W}_{0}$ such that it contains no information before editing. The final output of the edited layer combines the output of the original layer and the residual memory layer.
@@ 3.1 Editing: distributing knowledge across memory
Prior parametric approaches on knowledge editing Meng et al. (2023); Fang et al. (2025) typically store the knowledge of each edit across the entire trainable parameter space. However, unconstrained updates on all trainable parameters lead to severe catastrophic forgetting, as new edits quickly overwrite knowledge from previous ones, as shown in Figure 1. In this work, we draw motivation from the continual learning literature leveraging sparsity Mallya and Lazebnik (2018); Wortsman et al. (2020); Mirzadeh et al. (2020); Saha et al. (2021); Wang et al. (2022); Abbasi et al. (2022) to mitigate catastrophic forgetting, and propose a structured sparsification of input activations to the residual memory during the forward pass. This ensures that each edit modifies only a subset of the trainable parameters, promoting less overlap in representations and leading to more localized updates across different edits.
@@
A fundamental aspect of this sparse masking is the selection mechanism for the active indices. Intuitively, this mechanism should satisfy two different criteria: selection of feature diversity and semantic coherence. Diversity in feature selection encourages the model to consider a broader set of features, reducing the risk of overfitting on dominant features and mitigating catastrophic forgetting by distributing updates across the trainable parameters of the memory module. Semantic coherence ensures that semantically similar input prompts yield similar selected masks, facilitating the activation of relevant knowledge in the memory module for previously unseen yet semantically related queries. Therefore, we opt for structured sparsity Meier et al. (2008).
@@
defeating the purpose of structured and reusable knowledge storage. To address this limitation, we turn to a more robust fingerprinting mechanism based on the inherent structure of the activations themselves. For an input $\bm{a}=\bm{a}(\bm{x})\in\mathbb{R}^{D}$, we compute a top-k selection based on feature magnitudes, mapping $\bm{a}$ to a binary mask $\mathcal{T}(\bm{a})$. Denoting as $a_{(k)}$ the $k$-th largest value using order statistics notation, the mask is formally defined as $\mathcal{T}(\bm{a})_{j}=\mathds{1}[a_{j}\geq a_{(k)}],\;j\in[D]$. This strategy ensures that semantically similar input prompts—naturally producing similar activation patterns—yield similar masks, directly addressing the generalization shortcoming of random seeds. Given the structured and highly correlated nature of activations in LLMs, this step effectively acts as a Locality-Sensitive Hashing (LSH) Indyk and Motwani (1998) mechanism, mapping semantically similar input prompts to identical or closely related binary codes Norouzi et al. (2012).
@@ 3.2 Inference: activating knowledge conditional to inference prompt
During editing, we construct a database of sparse masks of each edited sample. Let $\bm{x}\in\mathcal{D}_{\textrm{edit}}$ be an edited prompt. Given that the layers prior to the introduced memory module are frozen, the application of the TopHash mechanism will yield the same sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ during both editing and inference, allowing the reliable identification of the prompt as part of the edit set by finding the exact match in the above database. Since transformers cluster semantically similar input prompts together due to their ability to generalize across paraphrases Reimers and Gurevych (2019), we expect that $\bm{x}$ should have activations, as well as TopHash masks, with high similarity w.r.t. its semantically rephrased variants. Conversely, semantically unrelated prompts will exhibit low similarity in both activation and mask space to the edited prompt $\bm{x}$.
We can compare the semantic similarity among prompts via the Hamming distance, denoted as $d_{H}$, between their respective masks Norouzi et al. (2012). By populating the database during editing, we can identify the semantically closest edited prompt to the prompt $\bm{x}$:
@@
The overlapping ratio $R_{\textrm{match}}$ measures the semantic relevance between a prompt and its closest edited prompt, enabling us to distinguish between different prompt types. Figure 4 presents the distribution of $R_{\textrm{match}}$ for edited prompts, rephrased variants, and irrelevant prompts across two benchmarks. We observe that these prompt categories exhibit distinct modes within the distribution, demonstrating the effectiveness of the proposed TopHash mechanism as a Locality-Sensitive Hashing (LSH) function Indyk and Motwani (1998) that clusters semantically similar input prompts.
@@
Prior approaches often require access to large corpora of irrelevant samples to enhance locality; either by sampling a different irrelevant sample during each edit (Wang et al., 2024) or by performing a forward pass on a corpora of up to 100,000100{,}000 samples to regularize editing (Meng et al., 2023; Fang et al., 2025). MEMOIR, on the other hand, eliminates the need for such large-scale irrelevant samples during editing and offers a more flexible alternative, as it relies on the introduced conditional knowledge activation mechanism in Equation 4 to decide whether to use the edited or pre-trained knowledge.